OpenStudy (anonymous):

Molly can prepare a report in 6 hours. Jere can prepare a report in only 4 hours. Which of the following can be used to determine the amount of time it would take for Molly and Jere to prepare a report together?

OpenStudy (anonymous):

1 over 6 plus 1 over x equals 1 over 4 1 over x plus 1 over 4 equals 1 over 6 1 over 4 plus 1 over 6 equals x over 10 1 over 6 plus 1 over 4 equals 1 over x

jimthompson5910 (jim_thompson5910):

Hint M = time it takes for molly to prepare the report (with no help from others) J = time it takes for Jere to prepare the report (with no help from others) \[\large \frac{1}{\text{Molly's Time}}+\frac{1}{\text{Jere's Time}} = \frac{1}{\text{Time if they work together}}\] \[\large \frac{1}{M}+\frac{1}{J} = \frac{1}{\text{Time if they work together}}\] I'll let you finish

OpenStudy (anonymous):

So would it be D?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Yay:) Can I ask another?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Ok hold on please

OpenStudy (anonymous):

Nathaniel can weld a railing in 75 minutes. Brenda can weld a railing 25 minutes faster. If they work together, how many minutes does it take them to weld the railing?

OpenStudy (anonymous):

25 minutes 30 minutes 50 minutes 150 minutes

jimthompson5910 (jim_thompson5910):

same idea, just different numbers

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{75} + \frac{1}{25} = \frac{1}{x}\] x = time in minutes it takes them to do the job together

OpenStudy (anonymous):

sO 50 minutes ?

jimthompson5910 (jim_thompson5910):

oh wait, I have the wrong equation

OpenStudy (anonymous):

ok:)

jimthompson5910 (jim_thompson5910):

`Nathaniel can weld a railing in 75 minutes. Brenda can weld a railing 25 minutes faster.` so Brenda takes 75-25 = 50 minutes \[\Large \frac{1}{75} + \frac{1}{50} = \frac{1}{x}\] isolate x

OpenStudy (anonymous):

how do I do that?

jimthompson5910 (jim_thompson5910):

what does 1/75 + 1/50 turn into?

jimthompson5910 (jim_thompson5910):

add them up like normal fractions

OpenStudy (anonymous):

150 is the gcf

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

well LCD actually but yeah

OpenStudy (anonymous):

2x+3x/150x=150/150x

OpenStudy (anonymous):

So 30 min?

jimthompson5910 (jim_thompson5910):

yes x = 30

OpenStudy (anonymous):

Yayyy!!!! Thank-you!

OpenStudy (anonymous):

Up for another?

jimthompson5910 (jim_thompson5910):

I'll help with one last one

OpenStudy (anonymous):

It takes Harland 15 minutes to rake all leaves in the front yard. If Trudy scatters the leaves while Harland rakes, it takes him 20 minutes to rake the leaves. How many minutes does it take Trudy to scatter all the leaves?

OpenStudy (anonymous):

Deal:)

OpenStudy (anonymous):

5 minutes 12 minutes 45 minutes 60 minutes

OpenStudy (anonymous):

I'm thinking 60 minutes

jimthompson5910 (jim_thompson5910):

what does 1/20 + 1/15 add up to?

OpenStudy (anonymous):

Wait! NVM! I did 1/15-1/20 instead of addition

OpenStudy (anonymous):

Um the LCD is 60

OpenStudy (anonymous):

4x/60x +3x/6x =60/60x

jimthompson5910 (jim_thompson5910):

oh I have a typo, let me fix it

OpenStudy (anonymous):

kk

jimthompson5910 (jim_thompson5910):

the 1/y should have been negative

OpenStudy (anonymous):

Ok :)

jimthompson5910 (jim_thompson5910):

sorry another typo, but I'm sure this time it's correct x = time it takes for Harland to rake all the leaves (without Trudy in the way) y = time it takes Trudy to scatter all the leaves (without Harland in the way) z = time it takes Harland to rake all the leaves with Trudy scattering the leaves at the same time \[\Large \frac{1}{x} - \frac{1}{y} = \frac{1}{z}\] \[\Large \frac{1}{15} - \frac{1}{y} = \frac{1}{20}\] \[\Large \frac{1}{15} - \frac{1}{y} + \frac{1}{y} = \frac{1}{20} + \frac{1}{y}\] \[\Large \frac{1}{15} = \frac{1}{20} + \frac{1}{y}\] \[\Large \frac{1}{15} - \frac{1}{20} = \frac{1}{20} + \frac{1}{y} - \frac{1}{20}\] \[\Large \frac{1}{15} - \frac{1}{20} = \frac{1}{y}\]

jimthompson5910 (jim_thompson5910):

not sure why I'm being so careless lol

OpenStudy (anonymous):

so would the answer be 60?

OpenStudy (anonymous):

It's fine<3

jimthompson5910 (jim_thompson5910):

anyways, you had the right idea in saying 1/15 - 1/20 that leads to 1/60 so y = 60

OpenStudy (anonymous):

Thanks so much for all your help:)))

jimthompson5910 (jim_thompson5910):

no problem